OpenStudy (anonymous):

integrate sqrt{(x^3-3)/x^11}

hartnn (hartnn):

\(\int \sqrt{\dfrac{x^3-3}{x^8x^3}}dx =\int \dfrac{1}{x^4} \sqrt{\dfrac{x^3-3}{x^3}}dx\) did you get this step ??

OpenStudy (anonymous):

yes

hartnn (hartnn):

then let u = \(\dfrac{x^3-3}{x^3}=1-3x^{-3}\)

hartnn (hartnn):

du=.... ?

OpenStudy (anonymous):

thanks :)

hartnn (hartnn):

so, can you solve further ?

OpenStudy (anonymous):

solved thanks again

hartnn (hartnn):

welcome ^_^